A valószínűségszámítás és matematikai statisztika dolgozatok eredménye

Megszerezte-e a jogot az aláírásra?

Agócs László Péter  igen
Arany Gábor  nem
Bálint Levente  nem
Csipkés László  nem
Eszenyi László  igen
Ferenczik Zoltán  nem
Fogarasi Edina  igen
Gerják Balázs  igen
Juhász Tibor  igen
Kis Bertalan  igen
Kocsis Gergely  igen
Kovács György  igen
Lázár Gábor  igen
Nagy Csaba  igen
Paluska Mihály  nem
Pári Viktor  nem
Pelejtei Gergely  nem
Sipos Nándor  igen
Szakácsi János  igen
Szalontai Gábor  igen
Szalontai József  igen
Tóth János  igen
Tóth Tamás  igen
Virág Tamás  igen
Zabos Katalin  igen

A valószínűségszámítás és matematikai statisztika dolgozatok eredménye

Megszerezte-e a jogot az aláírásra?

Dancsi József  nem
Emődi Szabolcs Árpád  nem
Fülöp Arnold  nem
Hamza Péter  nem
Hank Sándor  nem
Lörinczi József  igen
Molnár Angéla  igen
Molnár Tibor  igen
Nagy Gábor  igen
Nagy Zsolt  nem
Páli Gábor János  nem
Paper Andrea  nem
Papp Gyula  nem
Szabó Zsolt  igen
Szevelényi Attila  nem
Tilki Csaba  nem
Tordai Jácint  nem